OpenStudy (anonymous):

is -7 is solution to the -16+2=x+7 find the solution

OpenStudy (karatechopper):

-21

OpenStudy (anonymous):

ok..what?

OpenStudy (saifoo.khan):

-16 + 2 = -7 + 7 -14 = 0 Since they aren't same so it is not.

OpenStudy (anonymous):

x = -21